# Learned frequency-domain scattered wavefield solutions using neural operators

### v2

@@
Abstract
Solving the wave equation is essential to seismic imaging and inversion. The numerical solution of the Helmholtz equation, fundamental to this process, often encounters significant computational and memory challenges. We propose an innovative frequency-domain scattered wavefield modeling method employing neural operators adaptable to diverse seismic velocities. The source location and frequency information are embedded within the input background wavefield, enhancing the neural operator’s ability to process source configurations effectively. In addition, we utilize a single reference frequency, which enables scaling from larger-domain forward modeling to higher-frequency scenarios, thereby improving our method’s accuracy and generalization capabilities for larger-domain applications. Several tests on the OpenFWI datasets and realistic velocity models validate the accuracy and efficacy of our method as a surrogate model, demonstrating its potential to address the computational and memory limitations of numerical methods.
Index Terms:
@@ Abstract
Frequency-domain modeling, Fourier neural operators, scattered wavefield, neural PDE solvers
I Introduction
@@
Solving wave equations is pivotal for various geophysical applications like reverse time migration and full waveform inversion. However, the computational and memory bottlenecks often hinder real-time applications like microseismic imaging for CO2 monitoring. Frequency-domain simulation [1] offers a reduced dimension solution space for multi-scale waveform inversion. Yet, it still struggles with exponentially increasing memory demands with increasing model size and the necessity for repeated simulations across various frequencies.
@@
With the development of modern machine learning techniques comes the possibility of using neural networks to surrogate numerical modeling. Physics-informed neural networks (PINNs) have recently been employed for time-domain and frequency-domain wavefield modeling, demonstrating their potential in solving complex wave equations in irregular geometry through meshless solutions [2, 3, 4, 5, 6, 7, 8]. Despite their multi-source and multi-frequency capabilities [9], PINNs’ need for retraining or transfer learning for new velocities limits their generalization. Consequently, developing neural networks for a range of parametric PDEs becomes crucial. Neural operators [10] have shown great potential in solving for parametric PDEs, which provide instant PDE solutions given various coefficients and boundary conditions. Recently, they have also been used for seismic wavefield simulation, e.g., 2D/3D time-domain wavefield simulation [11, 12, 13]. However, learning seismic waveform modeling in the time domain faces several challenges: (1) Learning the mapping directly from input 2D velocities and source locations to 3D wavefield volumes (spanning 2D spatial and 1D time-lapse domains) requires huge memory cost [14]; (2) Learning the mapping between the snapshots at different times within recurrent composition though decreases the memory cost, will be subject to error propagation [15] as well as the relatively huge inference cost due to the iterative nature. Thus, learning in the frequency domain using neural operators, which avoids the above limitations, could efficiently handle parametric wave equations.
@@
Although current neural operator implementations for the frequency domain wavefield solutions offer memory efficiency, they typically often use the end-to-end framework to learn the mapping from the velocities, source locations, and frequencies to the corresponding wavefield solutions without additional feature engineering [14, 16, 17]. Specifically, they treat the source locations as binary masks and frequencies as an additional channel for inputting neural operators with constant values. However, without considering the physics of seismic wavefields, this approach offers suboptimal inputs for convolution-based feature extraction networks. Thus, it requires additional modifications to the network architecture, e.g., paralleled Fourier neural operator [17] or limits the target frequencies of modeling [14] to make the neural operators handle various velocities and source configurations (e.g., locations and frequency) simultaneously.
@@
To tackle this challenge, we propose learned frequency-domain scattered wavefield solutions using the Fourier neural operator. Unlike conventional methods that add source location and frequency as input channels, our proposed method implicitly embeds such information into an analytically evaluated background wavefield for a homogeneous background velocity. So, the neural operator focuses on mapping this background wavefield with the frequency and source information embedded in the scattered wavefield (or the full wavefield) for a given velocity model also as an input. The proposed method allows the neural operators to simultaneously handle the modeling for various velocities, frequencies, and source locations. Moreover, to improve generalization for larger-domain velocities beyond the training sample scope, we propose incorporating the single reference frequency [18] to enhance the generalization ability. We demonstrate the effectiveness of the proposed method on the OpenFWI dataset [19] and more realistic models, e.g., extracted from Overthrust models.
@@
Figure 1: The pipeline of the proposed method, where the labels denote the reference results using the finite-difference method, N denotes the number of samples, and H and W are the height and width of the images. The input has three channels, including velocity and real and imaginary parts of the analytical background wavefield, while the output consists of two channels representing real and imaginary parts of the scattered wavefield. For the FNO description, we refer the reader to [10].
@@
II Methodology
II-A Operator learning for Partial differential equations
@@
II-B Embedding the source configurations into the background wavefield
With $f_{\theta}(\mathbf{v},x_{s},z_{s},\omega)$ and reference pre-generated modeling results as labels, we can train the neural network $f$ to act as a surrogate modeling tool. However, this approach integrates the source location and the frequency as additional input channels, with a scalar value representing frequency and a binary mask denoting the source location [11]. While structurally convenient, such representations offer limited information for effectively utilizing learned kernels in feature extraction. Instead, here, we propose embedding such information, including the frequency and the source location, in the background wavefield $\mathbf{U}_{0}$, defined, for 2D media, as:
@@ II-B Embedding the source configurations into the background wavefield
II-C Learning for the scattered wavefield
As shown in many previous works [20, 15], learning the residual of the physical wavefield rather than directly learning the physical field itself will improve the accuracy of the neural network-based simulation. For example, Alkhalifah et al., (2020)[21] showed that solving for the scattered wavefield instead of directly learning wave equation solutions can help us avoid the point-source singularity. Inspired by those facts and considering we have a background wavefield with embedded frequency and source location information, we design the neural operator to predict the scattered wavefield, $\delta\mathbf{U}=\mathbf{U}-\mathbf{U}_{0}$, satisfying
@@ II-C Learning for the scattered wavefield
where $N$ is the number of training samples. As for the choice of the neural network, among many neural networks for parametric PDEs such as ResNet [22], Unet [23], UNO [24], LordNet [15], graph neural networks [25], transformers [26] and so on, we select the commonly used benchmark network, the Fourier neural operator (FNO) [10], to test the framework. The full pipeline of the approach is shown in Figure 1.
III Numerical Examples
We conducted tests on two velocity datasets, OpenFWI [19] and more realistic models, utilizing the Fourier Neural Operator (FNO) with four blocks (Fourier layer shown in Figure 1). Each block was configured with a truncated Fourier mode set to 48 and a Fourier layer width of 128. The activation function used here is the Gaussian Error Linear Unit (GELU) [27]. We employed an Adam optimizer to train the FNO, using a learning rate 1e-3. The training was conducted over 1000 epochs, with a batch size of 128 for the OpenFWI test and 64 for realistic models. The chosen frequency range, from 3 Hz to 21 Hz, aligns with the primary frequency band often used in seismic applications.
III-A Synthetic tests on OpenFWI dataset
Here, we first applied the proposed method to the OpenFWI dataset, focusing on the ’CurveVelA’ velocity class [19], as illustrated in Figure 3. This class of velocities includes curved layers, and the value of velocity within the layers gradually increases with depth. This test involved a comprehensive training set comprising 9,000 samples, encompassing various velocities, source locations, and frequencies, and a corresponding validating set comprising 1,000 samples. The validation error is estimated using the relative L2 norm error, which equals the Euclidean distance from the prediction to the ground truth divided over the Euclidean norm of the ground truth. Although the Fourier neural operator can learn the mapping between infinite function space, the training data here are discretized samples with a spatial resolution of 139$\times$139. The training curve and validation error analysis are shown in Figure 2, and as we can see, the training converges well.
@@ III-A Synthetic tests on OpenFWI dataset
Figure 4: The results of the network predictions for velocity models (Column 1) from the test set. The plots include the background wavefield (Column 2), the reference results from numerical solutions (Column 3), the predictions (Column 4), and the difference between the predictions and the numerical solutions (Column 5). In rows 1 and 3, we show the real part of the wavefields, while in rows 2 and 4, we show the imaginary part of the wavefields. The first two rows are the results of the same velocity and same frequency of 4.25 Hz but with different source locations. The third row is the result of the prediction on another velocity but with a frequency of 7.55 Hz, while the fourth row is the result of 9.88 Hz.
Figure 5: Samples of realistic velocity models used in the training.
@@
After the training, we evaluated the FNO for various unseen velocities, shown in Figure 4. For comparison, we use the optimal 9-point finite difference method to calculate the reference results with a spatial interval of 0.0125 km at both $x$ and $z$ directions. The predictions of FNO match well with the reference wavefields from numerical finite-difference modeling. It demonstrates that the network managed to extract the source location information and the frequency from the background wavefield and learn the relations between the various velocities and simulated complex wavefields.
III-B Comparison with conventional neural operator-based methods
@@
As mentioned above, conventional neural operator-based methods treat the source locations as binary masks and frequencies as additional channels. This approach offers suboptimal inputs for a convolution-based feature extraction network, yielding a degraded performance. In this subsection, we demonstrate this weakness of the conventional methods, especially compared to our proposed method for enhancing the wavefield representation. For a fair comparison, we use the same training and validation datasets, as well as the same configurations for the FNOs, to train the neural operator using the different approaches (pipelines). We specifically use three different approaches here: the conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), the conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), and our method. To have the evaluation metrics consistent and comparable, for the full wavefield output, the calculation of the relative L2 norm error is also done by the calculation of the L2 norm of the errors divided by the L2 norm of the scattered wavefield. The training loss curves and the validation error curves are shown in Figure 6. It is obvious that our approach, implicitly embedding the source location and frequency information into the background wavefield, can improve the convergence, as well as improve the accuracy and generalization ability a lot compared to conventional neural operator-based methods.
@@
We further compare their performance on the unseen velocities with 200 samples. As shown in Table I, our method achieves the best accuracy compared to the conventional approaches with almost the same computational cost.
TABLE I: The training cost and accuracy comparison between different approaches
@@
III-C More realistic velocity models
To further test the effectiveness of the proposed method, we choose more realistic velocity models to test, e.g., the combinations of Overthrust, Marmousi models, and so on. Figure 5 shows samples of velocities in the training dataset. The number of training samples is 5,400, while that for validation is 600. Figure 7 shows the training loss curve and the validation error at each epoch. We also train FNO up to 1000 epochs, and it converges well. We similarly obtain the numerical reference results on unseen cases with a spatial interval of 0.025 km in both $x$ and $z$ directions and compare them with the predictions from the FNO. Figure 8 shows a comparison between the numerical solutions and the network predictions. Like the OpenFWI dataset tests, the FNO predicts accurate wavefields instantly.
@@ III-C More realistic velocity models
Figure 8: The results of the network predictions for four realistic velocity models in which the order of the plots are the same as in Figure 4. The frequencies for each forward modeling are, 3.93 Hz, 5.00 Hz, 12.60 Hz, and 20.66 Hz from top to bottom.
III-D Generalization test with reference frequency:

### v3

@@
Learned frequency-domain scattered wavefield solutions using neural operators
Xinquan Huang1 and Tariq Alkhalifah1
1 Physical Science and Engineering division
Now at: University of Pennsylvania, Philadelphia, PA 19104, USA.
   King Abdullah University of Science and Technology
Thuwal 23955-6900
   Saudi Arabia
keywords:
Frequency-domain modeling; Neural operators; Scattered wavefield, Neural PDE solvers; Wave propagation
{summary}
Solving the wave equation is essential to seismic imaging and inversion. The numerical solution of the Helmholtz equation, fundamental to this process, often encounters significant computational and memory challenges. We propose an innovative frequency-domain scattered wavefield modeling method employing neural operators adaptable to diverse seismic velocities. The source location and frequency information are embedded within the input background wavefield, enhancing the neural operator’s ability to process source configurations effectively. In addition, we utilize a single reference frequency strategy, which enables scaling from larger-domain forward modeling to higher-frequency scenarios, thereby improving our method’s accuracy and generalization capabilities for larger-domain applications. Several tests on the OpenFWI datasets and realistic velocity models validate the accuracy and efficacy of our method as a surrogate model, demonstrating its potential to address the computational and memory limitations of numerical methods.
1 Introduction
@@
Solving wave equations is pivotal for various geophysical applications like reverse time migration and full waveform inversion. However, the computational and memory bottlenecks often hinder real-time applications like microseismic imaging for CO2 monitoring. Frequency-domain simulation (Marfurt, 1984) offers a reduced dimension solution space for multi-scale waveform inversion. Yet, it still struggles with exponentially increasing memory demands with increasing model size and the necessity for repeated simulations across various frequencies.
@@
With the development of modern machine learning techniques comes the possibility of using neural networks to surrogate numerical modeling. Physics-informed neural networks (PINNs) have recently been employed for time-domain and frequency-domain wavefield modeling, demonstrating their potential in solving complex wave equations in irregular geometry through meshless solutions (Moseley et al., 2020; Voytan & Sen, 2020; Alkhalifah et al., 2021; Song et al., 2021; Huang & Alkhalifah, 2022a; Zheng & Wang, 2023; Sethi et al., 2023; Wu et al., 2023). Despite their multi-source and multi-frequency capabilities (Huang et al., 2022), PINNs’ need for retraining or transfer learning for new velocities limits their generalization. Consequently, developing neural networks for a range of parametric PDEs becomes crucial. Neural operators (Li et al., 2020) have shown great potential in solving for parametric PDEs, which provide instant PDE solutions given various coefficients and boundary conditions. Recently, they have also been used for seismic wavefield simulation, e.g., 2D/3D time-domain wavefield simulation (Yang et al., 2023; Lehmann et al., 2023; Zhang et al., 2023; Lehmann et al., 2024). However, learning seismic waveform modeling in the time domain faces several challenges: (1) Learning the mapping directly from input 2D velocities and source locations to 3D wavefield volumes (spanning 2D spatial and 1D time-lapse domains) requires huge memory cost (Zou et al., 2023); (2) Learning the mapping between the snapshots at different times within recurrent composition though decreases the memory cost, will be subject to error propagation (Huang et al., 2024) as well as the relatively huge inference cost due to the iterative nature. Thus, learning in the frequency domain using neural operators, which avoids the above limitations, could efficiently handle parametric wave equations.
@@
Although current neural operator implementations for the frequency domain wavefield solutions offer memory efficiency, they typically often use the end-to-end framework to learn the mapping from the velocities, source locations, and frequencies to the corresponding wavefield solutions without additional feature engineering (Zou et al., 2023; Zhang et al., 2023; Li et al., 2023). Specifically, they treat the source locations as binary masks and frequencies as an additional channel for inputting neural operators with constant values. However, without considering the physics of seismic wavefields, this approach offers suboptimal inputs for convolution-based feature extraction networks. Thus, it requires additional modifications to the network architecture, e.g., paralleled Fourier neural operator (Li et al., 2023) or limits the target frequencies of modeling (Zou et al., 2023) to make the neural operators handle various velocities and source configurations (e.g., locations and frequency) simultaneously.
@@
To tackle this challenge, we propose learned frequency-domain scattered wavefield solutions using the Fourier neural operator. Unlike conventional methods that add source location and frequency as input channels, our proposed method implicitly embeds such information into an analytically evaluated background wavefield for a homogeneous background velocity. So, the neural operator focuses on mapping this background wavefield, with the frequency and source information embedded in it, to the scattered wavefield (or the full wavefield) for a given velocity model also as an input. The proposed method allows the neural operators to simultaneously handle the modeling for various velocities, frequencies, and source locations. Moreover, to improve generalization for larger-domain velocities beyond the training sample scope, we propose incorporating the single reference frequency (Huang & Alkhalifah, 2022b) to enhance the generalization ability. We demonstrate the effectiveness of the proposed method on the OpenFWI dataset (Deng et al., 2022) and more realistic models, e.g., extracted from Overthrust models.
@@
Figure 1: The pipeline of the proposed method, where the labels denote the reference results using the finite-difference method, $N$ denotes the number of samples, and $H$ and $W$ are the height and width of the images. The input has three channels, including velocity and real and imaginary parts of the analytical background wavefield, while the output consists of two channels representing real and imaginary parts of the scattered wavefield. For the FNO description, we refer the readers to Li et al. (2020).
@@
2 Methodology
2.1 Operator learning for Partial differential equations
@@
2.2 Embedding the source configurations into the background wavefield
With $f_{\theta}(\mathbf{v},x_{s},z_{s},\omega)$ and reference pre-generated modeling results as labels, we can train the neural network $f$ to act as a surrogate modeling tool. However, this approach integrates the source location and the frequency as additional input channels, with a scalar value representing frequency and a binary mask denoting the source location (Yang et al., 2023). While structurally convenient, such representations offer limited information for effectively utilizing learned kernels in feature extraction. Instead, here, we propose embedding such information, including the frequency and the source location, in the background wavefield $\mathbf{U}_{0}$, defined, for 2D media, as:
@@ 2.2 Embedding the source configurations into the background wavefield
2.3 Learning for the scattered wavefield
As mentioned earlier, we can train the network to learn either the full wavefield or the residual wavefield (in our case, the scattered wavefield). There exists a general observation in neural operator-based fluid simulation that learning the residual of the physical wavefield rather than directly learning the physical field itself will improve the accuracy of the neural network-based simulation (Wandel et al., 2021; Huang et al., 2024). In addition, Alkhalifah et al. (2020) showed that solving for the scattered wavefield instead of directly learning wave equation solutions can help us avoid the point-source singularity during the training of physics-informed neural networks. In our case, this allows the neural operator training to focus on the relatively lower amplitude scattered wavefield. Then, the full wavefield can be obtained by simply adding the explicitly computed background wavefield. This can be thought of as a form of residual or skip connection that has demonstrated its benefits in machine learning training (Wandel et al., 2021). Inspired by those facts and potential improvements, and considering we have a background wavefield with embedded frequency and source location information, we design the neural operator to predict the scattered wavefield, $\delta\mathbf{U}=\mathbf{U}-\mathbf{U}_{0}$, satisfying
@@ 2.3 Learning for the scattered wavefield
where $N$ is the number of training samples. As for the choice of the neural network, we select the commonly used benchmark network, the Fourier neural operator (FNO) (Li et al., 2020), to test the framework. The full pipeline of the approach is shown in Figure 1. To further justify the use of the scattered wavefield as output for our implementation, we include an ablation study on various output types in the discussion (Section 4).
3 Numerical Examples
We conducted tests on two velocity datasets, OpenFWI (Deng et al., 2022) and more realistic models, utilizing the Fourier Neural Operator (FNO) with four blocks (Fourier layer shown in Figure 1). Each block was configured with a truncated Fourier mode set to 48 and a Fourier layer width of 128. The activation function used here is the Gaussian Error Linear Unit (GELU) (Hendrycks & Gimpel, 2016). We do not normalize the datasets like what is usually done in computer vision tasks. Instead, we directly feed the velocities in km/s and the background wavefield into the neural network. Hence, for different input pairs, they can be used without any further processing. We employed an Adam optimizer to train the FNO, using a learning rate 1e-3. The training was conducted over 1000 epochs, with a batch size of 128 for the OpenFWI test and 64 for realistic models. The chosen frequency range, from 3 Hz to 21 Hz, aligns with the primary frequency band often used in seismic applications.
3.1 Synthetic tests on OpenFWI dataset
Here, we first applied the proposed method to the OpenFWI dataset, focusing on the ’CurveVelA’ velocity class (Deng et al., 2022), as illustrated in Figure 3. This class of velocities includes curved layers, and the value of velocity within the layers gradually increases with depth. This test involved a comprehensive training set comprising 9,000 samples, encompassing various velocities, source locations, and frequencies, and a corresponding validating set comprising 1,000 samples. For each sample, we randomly pick the velocity model from ’CurveVelA’, randomly set the source location on the surface, randomly draw a frequency value from 3Hz to 15 Hz, and choose the background velocity of 1.5 km/s to generate the corresponding background wavefield and scattered wavefield. The validation error is estimated using the relative L2 norm error, which equals the Euclidean distance from the prediction to the ground truth divided over the Euclidean norm of the ground truth. Although the Fourier neural operator can learn the mapping between infinite function space, the training data here are discretized samples with a spatial resolution of 139$\times$139. The training curve and validation error analysis are shown in Figure 2, and as we can see, the training converges well.
@@ 3.1 Synthetic tests on OpenFWI dataset
Figure 4: The results of the network predictions for unseen velocity models (Column 1) from the test set. The plots include the background wavefield (Column 2), the reference scattered wavefields from numerical solutions (Column 3), the predicted scattered wavefields (Column 4), and the difference between the predictions and the numerical solutions (Column 5). In rows 1 and 3, we show the real part of the wavefields, while in rows 2 and 4, we show the imaginary part of the wavefields. The first two rows are the results of the same velocity and same frequency of 4.25 Hz but with different source locations. The third row is the result of the prediction on another velocity but with a frequency of 7.55 Hz, while the fourth row is the result of 9.88 Hz.
@@
After the training, we evaluated the FNO for various unseen velocities, shown in Figure 4. The frequency is randomly drawn for each case from the frequency range of 3 Hz to 15 Hz, and the source location is set on the surface and randomly drawn from the range given by our surface dimension. For comparison, we use the optimal 9-point finite difference method to calculate the reference results with a spatial interval of 0.0125 km at both $x$ and $z$ directions. The predictions of FNO match well with the reference scattered wavefields from numerical finite-difference modeling. It demonstrates that the network managed to extract the source location information and the frequency from the background wavefield and learn the relations between the various velocities and simulated complex wavefields.
3.2 Comparison with conventional neural operator-based methods
@@
As mentioned above, conventional neural operator-based methods treat the source locations as binary masks and frequencies as additional channels. This approach offers suboptimal inputs for a convolution-based feature extraction network, yielding a degraded performance. In this subsection, we demonstrate the weakness of conventional methods, especially when compared to our proposed method for enhancing wavefield representation. For a fair comparison, we use the same training and validation datasets, as well as the same configurations for the FNOs, to train the neural operator using the different approaches (pipelines). We specifically use three different approaches here: the conventional inputs with full wavefield outputs (marked as ”Conventional in full out”), the conventional inputs with scattered wavefield outputs (marked as ”Conventional in residual out”), and our method. To have the evaluation metrics consistent and comparable, for the full wavefield output, the calculation of the relative L2 norm error is also done by the calculation of the L2 norm of the errors divided by the L2 norm of the scattered wavefield. The training loss curves and the validation error curves are shown in Figure 5. It is obvious that our approach, implicitly embedding the source location and frequency information into the background wavefield, can improve the convergence, as well as improve the accuracy and generalization ability a lot compared to conventional neural operator-based methods.
@@
We further compare their performance on the unseen velocities with 200 samples. As shown in Table 1, our method achieves the best accuracy compared to the conventional approaches with almost the same computational cost.
Table 1: The training cost and accuracy comparison between different approaches
@@
3.3 More realistic velocity models
To further test the effectiveness of the proposed method, we choose more realistic velocity models to test, e.g., the combinations of Overthrust, Marmousi models, and so on. We extracted 256×256 patches from the complete velocity models and applied shifting and reversal operations to generate an augmented set of velocity models. Figure 6 shows samples of velocities in the training dataset. The number of training samples is 5,400, while that for validation is 600. For each sample, the frequency was chosen between 3 Hz and 21 Hz, the velocity was randomly selected from the dataset, the source location was randomly positioned on the surface, and a constant background velocity of 2.8 km/s was used to generate the corresponding background and scattered wavefields. Figure 7 shows the training loss curve and the validation error at each epoch. We also train FNO up to 1000 epochs, and it converges well. We similarly obtain the numerical reference results on unseen cases with a spatial interval of 0.025 km in both $x$ and $z$ directions and compare them with the predictions from the FNO. Realistic models are considerably more complex than the ‘CurveVelA’ models, and the available velocity model samples do not fully capture the complete distribution of realistic models. Consequently, the generalization in realistic models is much more challenging. To demonstrate the effectiveness of our approach on complex models, we evaluate it under two scenarios. In the first scenario, we assess the generalization with respect to frequency and source location by using a velocity model from the training set while employing source locations and frequency configurations that were not encountered during training for the corresponding velocity model. The results, shown in Figure 8, indicate accurate predictions. In the second scenario, we evaluate performance on unseen velocity models, with four representative examples presented in Figure 9. Although the predictions generally align with the reference numerical results thanks to the inclusion of similar velocity models in the training dataset, we observe that higher frequencies tend to introduce larger errors. This finding underscores that generalizing across different velocity models is more challenging than generalizing over variations in source and frequency configurations. We contend that a larger set of training samples covering the entire distribution of velocity models is necessary to enhance the accuracy of neural operator-based simulations for complex velocity models.
Figure 6: Samples of realistic velocity models used in the training.
@@ 3.3 More realistic velocity models
Figure 8: The results of the network predictions for four realistic velocity models in which the order of the plots are the same as in Figure 4. The frequencies for each forward modeling are 5.00 Hz, 3.93 Hz, 20.66 Hz, and 12.60 Hz from top to bottom.
Figure 9: The results of the network predictions for four unseen realistic velocity models in which the order of the plots are the same as in Figure 4. The frequencies for each forward modeling are 7.19 Hz, 3.32 Hz, 11.98 Hz, and 13.76 Hz from top to bottom.
3.4 Generalization test with reference frequency